Bromine-82 has a half-life of 36 hours. A sample containing Br-82 was found to have an activity of 1.4x10^5 disintegrations/min. How many grams of Br-82 were present in the sample? Assume that there were no other radioactive nuclides in the sample.


1
Expert's answer
2021-02-16T07:44:41-0500

Q163849

Bromine-82 has a half-life of 36 hours. A sample containing Br-82 was found to have an activity of 1.4x10^5 disintegrations/min. How many grams of Br-82 were present in the sample? Assume that there were no other radioactive nuclides in the sample.


Solution:


The rate of radioactive disintegration is directly proportional to the number of radioactive elements present at that time.


"A \u221d N."


where A is the number of disintegration per unit time

N is the number of radioactive nuclei present at the given moment.


"A = λ N; \space \space \space \space \space λ \space is \space the \space decay \space constant." ..............Equation 1


From the given half-life we can find the decay constant by using the formula


"t_{1/2} = \frac{ln2}{ λ } = \frac{0.693}{ λ}"


which can be rearranged as


"λ = \frac{0.693}{t_{1/2}}"



Step 1: To find the decay constant in units of min-1.

"t_{1/2} = 36\space hours * \frac{60\space minutes }{1\space hour } = 2160 \space minutes"


So the decay constant will be equal to


"λ = \frac{0.693}{2160\space minutes } = 0.0003208\space min^{-1} = 3.208 * 10^{-4} min^{-1}"



Step 2: To find the number of nuclei present at that time.


In question, we are given

Rate of disintegration, A = 1.4x105 disintegrations/min.

and we have calculated, λ = 3.208 * 10-4 min-1 .


plug this in equation 1 we have


"A = \u03bb N;"


1.4x105 disintegrations/min = 3.208 * 10-4 min-1 * N


divide both the side by 3.208 * 10-4 min-1 we have.


"\frac{1.4*10^{5} disintegrations/min}{3.208 * 10^{-4} min^{-1}} = \frac{ 3.208 * 10^{-4} min^{-1} * N }{3.208 * 10^{-4} min^{-1} }"


"N = \frac{1.4*10^{5} disintegrations/min}{3.208 * 10^{-4} min^{-1}} = 4.364 * 10^{8} atoms."



So there are 4.364 * 108 atoms of bromine-82 present at that moment in the given sample.


Step 3: Convert the number of atoms of Bromine to grams.


We will need the atomic mass of Bromine-82. We are not provided the atomic mass in the given problem.

You can assume that atomic mass equal to 82 g/mole.

I have taken the atomic mass of Br-82 from a source and it is equal to 81.9168g/mol.



"moles \space of \space Br(82) = 4.364 * 10^{8} Br(82) \space atoms * \frac{1mol }{6.022*10^{23} \space atoms \space of \space Br(82)}"


= 7.247 * 10-16 mole of Br.



"grams \space of \space Br(82) = 7.247 * 10^{-16} mole\space of\space Br * \frac{81.9168\space grams \space of Br }{1\space mol\space Br}"


= 5.934 * 10-14 grams of Bromine-82.


in 2 significant the answer will be 5.9 * 10-14 grams of Bromine-82

So there were 5.9 *10-14 grams of bromine-82 in the given sample.




Please let me know if you have not understood any part of the problem.

Thank you. :)